# Latent Factor Analysis in Short Panels

### v2

@@ E.6 Orthogonal transformations and maximal invariant statistic
For any symmetric matrix $S$ and orthogonal matrix $O$ in $\mathbb{R}^{m\times m}$, we have $vech(O^{-1}SO)=\mathscr{R}(O)vech(S)$, where $\mathscr{R}(O)=\frac{1}{2}A_{m}^{\prime}(O^{\prime}\otimes O^{\prime})A_{m}$ is an orthogonal matrix, and $A_{m}$ is the duplication matrix defined in Appendix B. Transformations $\mathscr{R}(O)$ with orthogonal $O$ have the structure of a group: (a) $\mathscr{R}(I_{m})=I_{\frac{1}{2}m(m+1)}$, (b) $\mathscr{R}(O_{1})\mathscr{R}(O_{2})=\mathscr{R}(O_{2}O_{1})$, and (c) $[\mathscr{R}(O)]^{-1}=\mathscr{R}(O^{-1})$.
@@
Under Assumptions 1 and A.5, (a) $vech(\widetilde{\hat{S}^{*}})=\mathscr{R}(O)vech(\hat{S}^{*})$, (b) $\widetilde{\boldsymbol{X}}=\mathscr{R}(O){\boldsymbol{X}}$, (c) $I_{p}-\widetilde{\boldsymbol{X}}(\widetilde{\boldsymbol{X}}^{\prime}\widetilde{\boldsymbol{X}})^{-1}\widetilde{\boldsymbol{X}}^{\prime}=\mathscr{R}(O)[I_{p}-\boldsymbol{X}(\boldsymbol{X}^{\prime}\boldsymbol{X})^{-1}\boldsymbol{X}^{\prime}]\mathscr{R}(O)^{-1}$, (d) $\widetilde{\boldsymbol{R}}=\boldsymbol{R}\mathscr{R}(O)^{-1}$, (e) $\widetilde{\boldsymbol{R}}(I_{p}-\widetilde{\boldsymbol{X}}(\widetilde{\boldsymbol{X}}^{\prime}\widetilde{\boldsymbol{X}})^{-1}\widetilde{\boldsymbol{X}}^{\prime})={\boldsymbol{R}}(I_{p}-{\boldsymbol{X}}({\boldsymbol{X}}^{\prime}{\boldsymbol{X}})^{-1}{\boldsymbol{X}}^{\prime})\mathscr{R}(O)^{-1}$.
@@ E.7 Proofs of Lemmas 4-10
Proof of Lemma 5: (a) From Assumption 2, we have $E[\bar{\varepsilon}]=0$ and $V[\bar{\varepsilon}]=V\left[\frac{1}{n}\sum_{i,k=1}^{n}s_{i,k}V_{\varepsilon}^{1/2}w_{k}\right]\break=V_{\varepsilon}^{1/2}\frac{1}{n^{2}}\sum_{i,j,k,l=1}^{n}s_{i,k}s_{j,l}E[w_{k}w_{l}^{\prime}]V_{\varepsilon}^{1/2}=(\frac{1}{n^{2}}\sum_{i,j}^{n}\sigma_{i,j})V_{\varepsilon}$ where the $s_{i,k}$ are the elements of $\Sigma^{1/2}$. Now, $\frac{1}{n^{2}}\sum_{i,j=1}^{n}\sigma_{i,j}\leq C\frac{1}{n^{2}}\sum_{m=1}^{J_{n}}b_{m,n}^{1+\delta}=O(n^{\delta-1}\sum_{m=1}^{J_{n}}B_{m,n}^{1+\delta})=O(n^{\delta-1}J_{n}^{1/2}(\sum_{m=1}^{J_{n}}B_{m,n}^{2(1+\delta)})^{1/2})\break=o(n^{-1}J_{n}^{1/2})=o(n^{-1/2})$ from the Cauchy-Schwarz inequality and Assumptions 2 (c) and (d). Part (a) follows. To prove part (b), we use $E[\frac{1}{n}\varepsilon\varepsilon^{\prime}]\rightarrow V_{\varepsilon}^{0}$ and $V[vech((V_{\varepsilon}^{0})^{-1/2}(\frac{1}{n}\varepsilon\varepsilon^{\prime})(V_{\varepsilon}^{0})^{-1/2})]=\frac{1}{n}\Omega_{n}$ from the proof of Lemma 1, and $\frac{1}{n}\Omega_{n}=o(1)$ by Assumption A.3. Finally, to show part (c), write $\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}\beta_{i}^{\prime}=(V_{\varepsilon}^{0})^{1/2}\frac{1}{n}\sum_{i,j=1}^{n}s_{i,j}w_{j}\beta_{i}^{\prime}$. Then, $E[\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}\beta_{i}^{\prime}]=0$ while the variance of $vec(\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}\beta_{i}^{\prime})$ vanishes asymptotically since $V[vec(\frac{1}{n}\sum_{i,j=1}^{n}s_{i,j}w_{j}\beta_{i}^{\prime})]=\frac{1}{n^{2}}\sum_{i,j,m,l=1}^{n}s_{i,j}s_{m,l}\break(\beta_{i}\beta_{l}^{\prime})\otimes E[w_{j}w_{m}^{\prime}]=\frac{1}{n^{2}}\sum_{i,l=1}^{n}\sigma_{i,l}(\beta_{i}\beta_{l}^{\prime})\otimes I_{T}=o(1)$ under Assumptions 2 and A.2.
@@
Proof of Lemma 6: From the arguments in the proof of Proposition 7 with $\Psi_{y}=0$, the solution of the FOC is such that $\Psi_{F,j}^{\epsilon}=(\Lambda_{j}^{0}-R_{j}^{0})\Psi_{V_{\varepsilon}}^{\epsilon}(V_{\varepsilon}^{0})^{-1}F_{j}$ for $j=1,...,k$, and $diag(M_{F_{0},V_{\varepsilon}^{0}}\Psi_{V_{\varepsilon}}^{\epsilon}M_{F_{0},V_{\varepsilon}^{0}}^{\prime})\break=0$. Since $\Psi_{V_{\varepsilon}}^{\epsilon}$ is diagonal, the latter equation yields $M_{F_{0},V_{\varepsilon}^{0}}^{\odot 2}diag(\Psi_{V_{\varepsilon}}^{\epsilon})=0$. Under condition (a) of Lemma 6, we get $\Psi_{V_{\varepsilon}}^{\epsilon}=0$, which in turn implies $\Psi_{F}^{\epsilon}=0$. Thus, condition (a) is sufficient for local identification. It is also necessary to get uniqueness of the solution $\Psi_{V_{\varepsilon}}^{\epsilon}=0$. Moreover, conditions (a) and (b) of Lemma 6 are equivalent as shown in Appendix E.3. Further, conditions (a) and (c) are equivalent since $\Phi^{\odot 2}=M_{F_{0},V_{\varepsilon}^{0}}^{\odot 2}(V_{\varepsilon}^{0})^{2}$. Finally, let us show that condition (d) of Lemma 6 is both sufficient and necessary for local identification. The FOC for the Lagrangian problem are $\frac{\partial L_{0}(\theta)}{\partial\theta}-\frac{\partial g(\theta)^{\prime}}{\partial\theta}\lambda_{L}=0$ and $g(\theta)=0$, where $\lambda_{L}$ is the Lagrange multiplier vector. By expanding at first-order around $\theta_{0}$ and $\lambda_{0}=0$, we get $H_{0}\left(\begin{array}[]{cc}\theta-\theta_{0}\\ \lambda\end{array}\right)=0$, where $H_{0}:=\left(\begin{array}[]{cc}J_{0}&A_{0}\\ A_{0}^{\prime}&0\end{array}\right)$, with $A_{0}=\frac{\partial g(\theta_{0})^{\prime}}{\partial\theta}$, is the bordered Hessian. The parameters are locally identified if, and only if, $H_{0}$ is invertible. The latter condition is equivalent to $B_{0}^{\prime}J_{0}B_{0}$ being invertible. [^44] Indeed, we can show $|H_{0}|=(-1)^{\frac{1}{2}k(k-1)}|A_{0}^{\prime}A_{0}||B_{0}^{\prime}J_{0}B_{0}|$ by using $J_{0}A_{0}=0$, where the latter equality follows since the criterion is invariant to rotations of the latent factors.
@@
Proof of Lemma 7: By Assumption 2, $vec(W_{n})=(I_{k}\otimes V_{\varepsilon}^{1/2})\frac{1}{\sqrt{n}}\sum_{m=1}^{J_{n}}x_{m,n}$ where the $x_{m,n}:=\sum_{i,j\in I_{m}}s_{i,j}(\beta_{i}\otimes w_{j})$ are independent across $m$. Now, we apply the Liapunov CLT to show $\frac{1}{\sqrt{n}}\sum_{m=1}^{J_{n}}x_{m,n}\Rightarrow N(0,Q_{\beta}\otimes I_{T})$. We have $E[x_{m,n}]=0$ and $E[x_{m,n}x_{m,n}^{\prime}]=\left(\sum_{i,j\in I_{m}}\sigma_{i,j}\beta_{i}\beta_{j}^{\prime}\right)\otimes I_{T}$ and, by Assumption A.8, $\Omega_{W,n}:=\frac{1}{n}\sum_{m=1}^{J_{n}}E[x_{m,n}x_{m,n}^{\prime}]$ converges to the positive definite matrix $Q_{\beta}\otimes I_{T}$. Let us now check the multivariate Liapunov condition $\|\Omega_{W,n}^{-1/2}\|^{4}\frac{1}{n^{2}}\sum_{m=1}^{J_{n}}E[\|x_{m,n}\|^{4}]=o(1)$. Since $\|\Omega_{W,n}^{-1/2}\|=O_{p}(1)$, it suffices to prove $\frac{1}{n^{2}}\sum_{m=1}^{J_{n}}E[(x_{m,n}^{p,t})^{4}]=o(1)$, for any $p=1,...,k$ and $t=1,...,T$, where $x_{m,n}^{p,t}:=\sum_{i,j\in I_{m}}s_{i,j}\beta_{i,p}w_{j,t}$. For this purpose, Assumptions A.1 and A.2 yield $E[(x_{m,n}^{p,t})^{4}]\leq C(\sum_{i,j\in I_{m}}\sigma_{i,j})^{2}$. Then, we get $\frac{1}{n^{2}}\sum_{m=1}^{J_{n}}E[(x_{m,n}^{p,t})^{4}]\leq C\frac{1}{n^{2}}\sum_{m=1}^{J_{n}}b_{m,n}^{2(1+\delta)}\leq Cn^{2\delta}\sum_{m=1}^{J_{n}}B_{m,n}^{2(1+\delta)}=o(1)$ by Assumptions 2 (c) and (d). Part (a) of Lemma 7 follows. Moreover, $E[vech(\zeta_{m,n})x_{m,n}^{\prime}]=0$ and the proof of Lemma 1 imply part (b).
@@
Proof of Lemma 8: From the proof of Proposition 2 we have $M_{\boldsymbol{X}}\Omega_{Z^{*}}M_{\boldsymbol{X}}=M_{\boldsymbol{X}}\boldsymbol{R}^{\prime}\Omega\boldsymbol{R}M_{\boldsymbol{X}}$, where $\Omega=D+\kappa I_{T(T+1)/2}=[\psi(0)-2q]D(0)+\sum_{h=1}^{T-1}\psi(h)[\tilde{D}(h)+\bar{D}(h)]+(q+\kappa)I_{T(T+1)/2}$. Then, since the columns of $\boldsymbol{R}$ are orthonormal, we get $M_{\boldsymbol{X}}\Omega_{Z^{*}}M_{\boldsymbol{X}}=[\psi(0)-2q]M_{\boldsymbol{X}}\boldsymbol{R}^{\prime}D(0)\boldsymbol{R}M_{\boldsymbol{X}}+\sum_{h=1}^{T-1}\psi(h)M_{\boldsymbol{X}}\boldsymbol{R}^{\prime}\tilde{D}(h)\boldsymbol{R}M_{\boldsymbol{X}}+\sum_{h=1}^{T-1}\psi(h)M_{\boldsymbol{X}}\boldsymbol{R}^{\prime}\bar{D}(h)\boldsymbol{R}M_{\boldsymbol{X}}+(q+\kappa)M_{\boldsymbol{X}}$. Now, we show that the the first two terms in this sum are nil. We have $G^{\prime}E_{t,t}G=Q^{\prime}V_{\varepsilon}^{1/2}E_{t,t}V_{\varepsilon}^{1/2}Q=V_{\varepsilon,tt}Q^{\prime}E_{t,t}Q$ and thus $vech(G^{\prime}E_{t,t}G)=V_{\varepsilon,tt}vech(Q^{\prime}E_{t,t}Q)=V_{\varepsilon,tt}\boldsymbol{R}^{\prime}vech(E_{t,t})$ (see the proof of Proposition 2). Hence, the kernel of matrix $M_{\boldsymbol{X}}$ is spanned by vectors $\boldsymbol{R}^{\prime}vech(E_{t,t})$, for $t=1,...,T$. We deduce that $M_{\boldsymbol{X}}\boldsymbol{R}^{\prime}D(0)=0$ and $M_{\boldsymbol{X}}\boldsymbol{R}^{\prime}\tilde{D}(h)\boldsymbol{R}M_{\boldsymbol{X}}=0$. Furthermore, from $I_{T(T+1)/2}=2\sum_{t=1}^{T}vech(E_{t,t})vech(E_{t,t})^{\prime}+\sum_{t<s}vech(E_{t,s}+E_{s,t})vech(E_{t,s}+E_{s,t})^{\prime}=2D(0)+\sum_{h=1}^{T-1}\bar{D}(h)$, we get $M_{\boldsymbol{X}}=M_{\boldsymbol{X}}\boldsymbol{R}^{\prime}I_{T(T+1)/2}\boldsymbol{R}M_{\boldsymbol{X}}=\sum_{h=1}^{T-1}M_{\boldsymbol{X}}\boldsymbol{R}^{\prime}\bar{D}(h)\boldsymbol{R}M_{\boldsymbol{X}}$. The conclusion follows.

### v3

@@ B.6 Orthogonal transformations and maximal invariant statistic
For any symmetric matrix $S$ and orthogonal matrix $O$ in $\mathbb{R}^{m\times m}$, we have $vech(O^{-1}SO)=\mathscr{R}(O)vech(S)$, where $\mathscr{R}(O)=\frac{1}{2}A_{m}^{\prime}(O^{\prime}\otimes O^{\prime})A_{m}$ is an orthogonal matrix, and $A_{m}$ is the duplication matrix defined in Appendix A. Transformations $\mathscr{R}(O)$ with orthogonal $O$ have the structure of a group: (a) $\mathscr{R}(I_{m})=I_{\frac{1}{2}m(m+1)}$, (b) $\mathscr{R}(O_{1})\mathscr{R}(O_{2})=\mathscr{R}(O_{2}O_{1})$, and (c) $\mathscr{R}(O)]^{-1}=\mathscr{R}(O^{-1})$.
@@
Under Assumptions 1 and 7, (a) $vech(\widetilde{\hat{S}^{*}})=\mathscr{R}(O)vech(\hat{S}^{*})$, (b) $\widetilde{\boldsymbol{X}}=\mathscr{R}(O){\boldsymbol{X}}$, (c) $I_{p}-\widetilde{\boldsymbol{X}}(\widetilde{\boldsymbol{X}}^{\prime}\widetilde{\boldsymbol{X}})^{-1}\widetilde{\boldsymbol{X}}^{\prime}=\mathscr{R}(O)[I_{p}-\boldsymbol{X}(\boldsymbol{X}^{\prime}\boldsymbol{X})^{-1}\boldsymbol{X}^{\prime}]\mathscr{R}(O)^{-1}$, (d) $\widetilde{\boldsymbol{R}}=\boldsymbol{R}\mathscr{R}(O)^{-1}$, (e) $\widetilde{\boldsymbol{R}}(I_{p}-\widetilde{\boldsymbol{X}}(\widetilde{\boldsymbol{X}}^{\prime}\widetilde{\boldsymbol{X}})^{-1}\widetilde{\boldsymbol{X}}^{\prime})={\boldsymbol{R}}(I_{p}-{\boldsymbol{X}}({\boldsymbol{X}}^{\prime}{\boldsymbol{X}})^{-1}{\boldsymbol{X}}^{\prime})\mathscr{R}(O)^{-1}$.
@@ B.7 Proofs of Lemmas 5-11
Proof of Lemma 6: (a) From Assumption 2, we have $E[\bar{\varepsilon}]=0$ and $V[\bar{\varepsilon}]=V\left[\frac{1}{n}\sum_{i,k=1}^{n}s_{i,k}V_{\varepsilon}^{1/2}w_{k}\right]\penalty-10000=V_{\varepsilon}^{1/2}\frac{1}{n^{2}}\sum_{i,j,k,l=1}^{n}s_{i,k}s_{j,l}E[w_{k}w_{l}^{\prime}]V_{\varepsilon}^{1/2}=(\frac{1}{n^{2}}\sum_{i,j}^{n}\sigma_{i,j})V_{\varepsilon}$ where the $s_{i,k}$ are the elements of $\Sigma^{1/2}$. Now, $\frac{1}{n^{2}}\sum_{i,j=1}^{n}\sigma_{i,j}\leq C\frac{1}{n^{2}}\sum_{m=1}^{J_{n}}b_{m,n}^{1+\delta}=O(n^{\delta-1}\sum_{m=1}^{J_{n}}B_{m,n}^{1+\delta})=O(n^{\delta-1}J_{n}^{1/2}(\sum_{m=1}^{J_{n}}B_{m,n}^{2(1+\delta)})^{1/2})\penalty-10000=o(n^{-1}J_{n}^{1/2})=o(n^{-1/2})$ from the Cauchy-Schwarz inequality and Assumptions 2 (c) and (d). Part (a) follows. To prove part (b), we use $E[\frac{1}{n}\varepsilon\varepsilon^{\prime}]\rightarrow V_{\varepsilon}^{0}$ and $V[vech((V_{\varepsilon}^{0})^{-1/2}(\frac{1}{n}\varepsilon\varepsilon^{\prime})(V_{\varepsilon}^{0})^{-1/2})]=\frac{1}{n}\Omega_{n}$ from the proof of Lemma 2, and $\frac{1}{n}\Omega_{n}=o(1)$ by Assumption 5. Finally, to show part (c), write $\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}\beta_{i}^{\prime}=(V_{\varepsilon}^{0})^{1/2}\frac{1}{n}\sum_{i,j=1}^{n}s_{i,j}w_{j}\beta_{i}^{\prime}$. Then, $E[\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}\beta_{i}^{\prime}]=0$ while the variance of $vec(\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}\beta_{i}^{\prime})$ vanishes asymptotically since $V[vec(\frac{1}{n}\sum_{i,j=1}^{n}s_{i,j}w_{j}\beta_{i}^{\prime})]=\frac{1}{n^{2}}\sum_{i,j,m,l=1}^{n}s_{i,j}s_{m,l}\penalty-10000(\beta_{i}\beta_{l}^{\prime})\otimes E[w_{j}w_{m}^{\prime}]=\frac{1}{n^{2}}\sum_{i,l=1}^{n}\sigma_{i,l}(\beta_{i}\beta_{l}^{\prime})\otimes I_{T}=o(1)$ under Assumptions 2 and 4.
@@
Proof of Lemma 7: From the arguments in the proof of Proposition 6 with $\Psi_{y}=0$, the solution of the FOC is such that $\Psi_{F,j}^{\epsilon}=(\Lambda_{j}^{0}-R_{j}^{0})\Psi_{V_{\varepsilon}}^{\epsilon}(V_{\varepsilon}^{0})^{-1}F_{j}$ for j=1,,kj=1,...,k, and $diag(M_{F_{0},V_{\varepsilon}^{0}}\Psi_{V_{\varepsilon}}^{\epsilon}M_{F_{0},V_{\varepsilon}^{0}}^{\prime})\penalty-10000=0$. Since $\Psi_{V_{\varepsilon}}^{\epsilon}$ is diagonal, the latter equation yields $M_{F_{0},V_{\varepsilon}^{0}}^{\odot 2}diag(\Psi_{V_{\varepsilon}}^{\epsilon})=0$. Under condition (a) of Lemma 7, we get $\Psi_{V_{\varepsilon}}^{\epsilon}=0$, which in turn implies $\Psi_{F}^{\epsilon}=0$. Thus, condition (a) is sufficient for local identification. It is also necessary to get uniqueness of the solution $\Psi_{V_{\varepsilon}}^{\epsilon}=0$. Moreover, conditions (a) and (b) of Lemma 7 are equivalent as shown in Appendix B.3. Further, conditions (a) and (c) are equivalent since $\Phi^{\odot 2}=M_{F_{0},V_{\varepsilon}^{0}}^{\odot 2}(V_{\varepsilon}^{0})^{2}$. Finally, let us show that condition (d) of Lemma 7 is both sufficient and necessary for local identification. The FOC for the Lagrangian problem are $\frac{\partial L_{0}(\theta)}{\partial\theta}-\frac{\partial g(\theta)^{\prime}}{\partial\theta}\lambda_{L}=0$ and $g(\theta)=0$, where $\lambda_{L}$ is the Lagrange multiplier vector. By expanding at first-order around $\theta_{0}$ and $\lambda_{0}=0$, we get $H_{0}\left(\begin{array}[]{cc}\theta-\theta_{0}\\ \lambda\end{array}\right)=0$, where $H_{0}:=\left(\begin{array}[]{cc}J_{0}&A_{0}\\ A_{0}^{\prime}&0\end{array}\right)$, with $A_{0}=\frac{\partial g(\theta_{0})^{\prime}}{\partial\theta}$, is the bordered Hessian. The parameters are locally identified if, and only if, $H_{0}$ is invertible. The latter condition is equivalent to $B_{0}^{\prime}J_{0}B_{0}$ being invertible. [^17] Indeed, we can show $|H_{0}|=(-1)^{\frac{1}{2}k(k-1)}|A_{0}^{\prime}A_{0}||B_{0}^{\prime}J_{0}B_{0}|$ by using $J_{0}A_{0}=0$, where the latter equality follows since the criterion is invariant to rotations of the latent factors.
@@
Proof of Lemma 8: By Assumption 2, $vec(W_{n})=(I_{k}\otimes V_{\varepsilon}^{1/2})\frac{1}{\sqrt{n}}\sum_{m=1}^{J_{n}}x_{m,n}$ where the $x_{m,n}:=\sum_{i,j\in I_{m}}s_{i,j}(\beta_{i}\otimes w_{j})$ are independent across $m$. Now, we apply the Liapunov CLT to show $\frac{1}{\sqrt{n}}\sum_{m=1}^{J_{n}}x_{m,n}\Rightarrow N(0,Q_{\beta}\otimes I_{T})$. We have $E[x_{m,n}]=0$ and $E[x_{m,n}x_{m,n}^{\prime}]=\left(\sum_{i,j\in I_{m}}\sigma_{i,j}\beta_{i}\beta_{j}^{\prime}\right)\otimes I_{T}$ and, by Assumption 10, $\Omega_{W,n}:=\frac{1}{n}\sum_{m=1}^{J_{n}}E[x_{m,n}x_{m,n}^{\prime}]$ converges to the positive definite matrix $Q_{\beta}\otimes I_{T}$. Let us now check the multivariate Liapunov condition $\|\Omega_{W,n}^{-1/2}\|^{4}\frac{1}{n^{2}}\sum_{m=1}^{J_{n}}E[\|x_{m,n}\|^{4}]=o(1)$. Since $\|\Omega_{W,n}^{-1/2}\|=O_{p}(1)$, it suffices to prove $\frac{1}{n^{2}}\sum_{m=1}^{J_{n}}E[(x_{m,n}^{p,t})^{4}]=o(1)$, for any p=1,,kp=1,...,k and t=1,,Tt=1,...,T, where $x_{m,n}^{p,t}:=\sum_{i,j\in I_{m}}s_{i,j}\beta_{i,p}w_{j,t}$. For this purpose, Assumptions 3 and 4 yield $E[(x_{m,n}^{p,t})^{4}]\leq C(\sum_{i,j\in I_{m}}\sigma_{i,j})^{2}$. Then, we get $\frac{1}{n^{2}}\sum_{m=1}^{J_{n}}E[(x_{m,n}^{p,t})^{4}]\leq C\frac{1}{n^{2}}\sum_{m=1}^{J_{n}}b_{m,n}^{2(1+\delta)}\leq Cn^{2\delta}\sum_{m=1}^{J_{n}}B_{m,n}^{2(1+\delta)}=o(1)$ by Assumptions 2 (c) and (d). Part (a) of Lemma 8 follows. Moreover, $E[vech(\zeta_{m,n})x_{m,n}^{\prime}]=0$ and the proof of Lemma 2 imply part (b).
@@
Proof of Lemma 9: From the proof of Proposition 1 we have $M_{\boldsymbol{X}}\Omega_{Z^{*}}M_{\boldsymbol{X}}=M_{\boldsymbol{X}}\boldsymbol{R}^{\prime}\Omega\boldsymbol{R}M_{\boldsymbol{X}}$, where $\Omega=D+\kappa I_{T(T+1)/2}=[\psi(0)-2q]D(0)+\sum_{h=1}^{T-1}\psi(h)[\tilde{D}(h)+\bar{D}(h)]+(q+\kappa)I_{T(T+1)/2}$. Then, since the columns of $\boldsymbol{R}$ are orthonormal, we get $M_{\boldsymbol{X}}\Omega_{Z^{*}}M_{\boldsymbol{X}}=[\psi(0)-2q]M_{\boldsymbol{X}}\boldsymbol{R}^{\prime}D(0)\boldsymbol{R}M_{\boldsymbol{X}}+\sum_{h=1}^{T-1}\psi(h)M_{\boldsymbol{X}}\boldsymbol{R}^{\prime}\tilde{D}(h)\boldsymbol{R}M_{\boldsymbol{X}}+\sum_{h=1}^{T-1}\psi(h)M_{\boldsymbol{X}}\boldsymbol{R}^{\prime}\bar{D}(h)\boldsymbol{R}M_{\boldsymbol{X}}+(q+\kappa)M_{\boldsymbol{X}}$. Now, we show that the the first two terms in this sum are nil. We have $G^{\prime}E_{t,t}G=Q^{\prime}V_{\varepsilon}^{1/2}E_{t,t}V_{\varepsilon}^{1/2}Q=V_{\varepsilon,tt}Q^{\prime}E_{t,t}Q$ and thus $vech(G^{\prime}E_{t,t}G)=V_{\varepsilon,tt}vech(Q^{\prime}E_{t,t}Q)=V_{\varepsilon,tt}\boldsymbol{R}^{\prime}vech(E_{t,t})$ (see the proof of Proposition 1). Hence, the kernel of matrix $M_{\boldsymbol{X}}$ is spanned by vectors $\boldsymbol{R}^{\prime}vech(E_{t,t})$, for t=1,,Tt=1,...,T. We deduce that $M_{\boldsymbol{X}}\boldsymbol{R}^{\prime}D(0)=0$ and $M_{\boldsymbol{X}}\boldsymbol{R}^{\prime}\tilde{D}(h)\boldsymbol{R}M_{\boldsymbol{X}}=0$. Furthermore, from $I_{T(T+1)/2}=2\sum_{t=1}^{T}vech(E_{t,t})vech(E_{t,t})^{\prime}+\sum_{t<s}vech(E_{t,s}+E_{s,t})vech(E_{t,s}+E_{s,t})^{\prime}=2D(0)+\sum_{h=1}^{T-1}\bar{D}(h)$, we get $M_{\boldsymbol{X}}=M_{\boldsymbol{X}}\boldsymbol{R}^{\prime}I_{T(T+1)/2}\boldsymbol{R}M_{\boldsymbol{X}}=\sum_{h=1}^{T-1}M_{\boldsymbol{X}}\boldsymbol{R}^{\prime}\bar{D}(h)\boldsymbol{R}M_{\boldsymbol{X}}$. The conclusion follows.